ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

1. Logic

1991 MS Classification: 3, 6 Burger, Isabella, Rand Afrikaans University, South Africa Boolean algebras and information-dependent entailment relations Consider a Lindenbaum-Tarski algebra B = (B, ⊧, ¬, ∧, ∨, ⊥, ⊤). We are interested in inducing new Boolean algebras by generalizing the classical entailment relation ⊧ to a data-dependent relation in light of information which is indeterminate, but for which we know the range of possible variation of its logical strength. The algebraic relationships between B and the induced Boolean algebras will be discussed as well as links with the work of Theo Kuipers.